OpenStudy (anonymous):

Determine, by graphing, whether the system of equations has one solution, no solution, or infinitely many solutions. If it has one solution, write it as an ordered pair. y=3x x+21=-2y a.infinite many b.one solution (-3,-9) c.no solution d.one solution (0,0)

OpenStudy (anonymous):

\[x+21=-2y\] Try to get y by itself first so you can compare the two equation.s

OpenStudy (anonymous):

Having trouble setting it up romero

OpenStudy (anonymous):

y=-x/2+21/2

OpenStudy (anonymous):

y=-x/2-21/2 Forgot a negative sign Anyways they want you to graph it.

OpenStudy (anonymous):

and i look for intersection

OpenStudy (anonymous):

Yes if there is one.

OpenStudy (anonymous):

its a straight line through like negative 4

OpenStudy (anonymous):

well it crosses x at negative

OpenStudy (anonymous):

So the solution has to be in the negative side. Which of the answers give you a negative point?

OpenStudy (anonymous):

B

OpenStudy (anonymous):

Yup b.one solution (-3,-9)

OpenStudy (anonymous):

k getting the hang of it ha

OpenStudy (anonymous):

Nice!! One thing I always like to check first is the slope of the two lines

OpenStudy (anonymous):

It can tell me right away if there is a solution or not. Remember same slope means parallel lines which means two things. No solution or infinite solution.